# Separate and Conquer: Decoupling Co-occurrence via Decomposition and Representation for Weakly Supervised Semantic Segmentation

### 29 Feb 2024 v2

@@ Abstract
Attributed to the frequent coupling of co-occurring objects and the limited supervision from image-level labels, the challenging co-occurrence problem is widely present and leads to false activation of objects in weakly supervised semantic segmentation (WSSS). In this work, we devise a ’Separate and Conquer’ scheme SeCo to tackle this issue from dimensions of image space and feature space. In the image space, we propose to ’separate’ the co-occurring objects with image decomposition by subdividing images into patches. Importantly, we assign each patch a category tag from Class Activation Maps (CAMs), which spatially helps remove the co-context bias and guide the subsequent representation. In the feature space, we propose to ’conquer’ the false activation by enhancing semantic representation with multi-granularity knowledge contrast. To this end, a dual-teacher-single-student architecture is designed and tag-guided contrast is conducted to guarantee the correctness of knowledge and further facilitate the discrepancy among co-occurring objects. We streamline the multi-staged WSSS pipeline end-to-end and tackle co-occurrence without external supervision. Extensive experiments are conducted, validating the efficiency of our method tackling co-occurrence and the superiority over previous single-staged and even multi-staged competitors on PASCAL VOC and MS COCO. Code will be available here.
@@ 1 Introduction
Figure 1: (a) Co-occurrence issue. Targets marked by star (horse and boat) are falsely activated. (b) To solve this issue, we propose a single-staged framework SeCo, which acts in ’separate and conquer’ manner that efficiently tackles co-occurrence issue without external supervision. It initially separates spatial con-texts in the image space and then conquer false activation in feature space. (c) The proposed SeCo accurately localises the co-occurring targets.
@@
Formally, the pipeline for WSSS with image-level labels consists of three steps, i.e., initially training a classification model to generate CAM seeds [44], then refining CAMs to generate pseudo labels, and finally training a segmentation network with the pseudo labels and taking it as the final model to inference [2, 47]. Such WSSS paradigm can be further divided into multi-staged and single-staged groups. For multi-staged WSSS methods [24, 46, 27], the classification and segmentation models need to be trained progressively. It intends to have better segmentation performance while more complicated at training streamline. The single-staged WSSS methods [2, 36, 37, 31] share the encoder for classification and segmentation networks, thus can be trained end-to-end. It is more efficient to optimize while holding inferior performance to the multi-staged. In this study, we focus on the most challenging WSSS paradigm with image-level annotation and streamline the paradigm end-to-end.
@@
For both single- and multi-staged WSSS, generating reliable CAMs from image-level labels is the first and fundamental step for the performance [19, 48]. However, since objects intend to co-occur together, such as {train, railroad}, {boat, water}, {horse, person}, etc., it is inevitable to tackle the co-occurrence of objectives in WSSS. The challenging co-occurrence problem is widespread and often leads to false activation [5, 22, 46, 20], as shown in Fig. 1 (a). Although most existing works have succeeded in completing CAMs, they neglect the existence of the co-occurring contexts, consequently limited to tackling the false activation and bottlenecked in WSSS performance. Recently, introducing external supervision or human prior is proposed to tackle the co-occurrence problem.  [46, 27] leverages CLIP model [33] to help distinguish between coupled contexts. [20] elaborately applies hard out-of-distribution samples to suppress spurious background cues. [39] constructs additional co-categories to tackle this issue. Although impressive, they heavily rely on external data or elaborate design to tackle co-occurrence, which impedes the real-world applications with complex relations among categories.
@@
Essentially, the co-occurrence problem arises because the co-occurring contexts coupled in images confuse networks and incur semantic bias during the feature representation, resulting in false positive pixels activated with high probability. Previous methods ignore the importance of separating co-contexts before representation, thus requiring external data or design to tackle co-occurrence. Based on the analysis above, we argue that separating the coupled objects initially to avoid noise and then enhancing category-specific representation provides a new insight to tackle this issue without external supervision.
@@
As illustrated in Fig. 1 (b), we propose a single-staged WSSS framework SeCo that does not require any extra supervision to tackle the problem. Our key insight lies in the ’separate and conquer’ training scheme that decouples co-occurrence in image space and feature space by image decomposition and representation enhancement, respectively. (1) To separate the spatial dependence of co-contexts in image space, we propose to decompose the integral image into multiple patches containing single category information. Previous patch-based method [17] simply giving image-level labels to patches, but the image labels cannot help differentiate the co-contexts at patch level (see Sec. 2.1). Instead, we further explore the strategy to spatially separate coupled objects and focus on tagging each patch. Specifically, patch-level category tags from CAMs are designed for each patch and help to identify the co-contexts at patch level. Since CAMs inevitably brings noise in tags, a similarity-based rectification method is designed to revise the noisy tags. In addition, a tag memory pool is constructed to store all history tags, guiding the subsequent patch representation. (2) To conquer the co-context confusion and enhance the semantic representation in feature space, we design a dual-teacher single-student architecture and promote discrepancy among co-categories. We first build a global teacher to extract category knowledge from integral images. The knowledge provides class centroids for the student in patch representation and helps to push apart co-contexts. Considering the trade-offs between the separation of co-categories and the destruction on global contexts, we share the encoders of both branches to provides complementary information for patch and image semantics. In addition, patch semantic knowledge is extracted from a patch-level semantic reservoir maintained by a local teacher. Guided by the tag memory pool, the knowledge helps reduce the noise during the representation and push apart co-contexts while pulling together those within the same category at a fine-grained level. (3) Along with the category tags and the extracted knowledge, multi-granularity contrast is further proposed across the whole dataset to decouple the co-contexts deeply.
@@
Extensive experiments are conducted on PASCAL VOC and MS COCO, validating its effectiveness on tackling co-occurrence (as shown in Fig. 1 (c)) and the superiority over previous single-staged and multi-staged competitors.
@@ 2 Related Works
Figure 2: (a) Architecture of the proposed SeCo to tackle co-occurrence issue. Specifically, integral images are firstly sent to the global teacher (G-Teacher) to extract the category knowledge and CAMs. Then category tags from CAM seeds are allocated to patches. With tags, two views of patches by different augmentations, i.e., weak data augmentation (W.T.) and strong augmentation (S.T.), are sent to the student and local teacher (L-Teacher) branch, respectively. The local teacher stores all the history patches and category tags and generalises the patch semantic knowledge. Finally, two contrastive losses, ${L}_{{LiG}}$ and ${L}_{{LiL}}$, are conducted to guarantee the decoupling. In addition, CAMs from global teacher are refined as pseudo labels to train the segmentation network. Since the encoders of segmentation model and classification model are shared, our WSSS pipeline can be trained end-to-end. (b) Illustrated essence of the key components in SeCo. More details are introduced in Sec. 3.2
@@ 2.1 Learning from Local Semantics
The image-level labels provide limited supervision to generate high-quality CAMs, which motivates many researches to dig complementary information from local semantics. Decomposing a whole image into local patches offers a practical implementation. L2G [17] crops patches to mine different views of semantics and generates more complete CAMs. PPL [23] utilizes feature patches to explore scattered local details and generate CAMs to cover the whole objects. ToCo [37] extracts local semantics from patches and activate more non-discriminative areas. However, simply giving image-level labels to each patch like L2G cannot identify co-categories at patch level, or grouping them into foregrounds and backgrounds as ToCo fails to tell co-categories in the foregrounds, which introduces much co-occurring noise during patch representation, thus suffering from false activation. To this end, we observe that rich semantics from CAMs can be potentially used to tag each patch, which helps differentiate co-contexts and guide category representation at a patch level. To the best of our knowledge, we are the first to report that assigning class-aware tags to patches effectively separates co-categories and dramatically reduce the noise during the representation.
@@ 2.2 Contrastive Learning & Knowledge Distillation
Contrastive learning (CL) [9, 15, 43] and knowledge distillation (KD) [16, 4] are two prevalent techniques to promote feature representation. Inspired by it, RCA [52] and ToCo [37] conduct region-level contrast to focus on non-discriminative pixels. PPC [13] leverage prototype-based [51] contrast to expand CAMs. Apart from incorporating CL, SCD [49] and L2G [17] introduce knowledge distillation into WSSS and succeed in completing CAMs. However, since the co-occurring pixels intend to be falsely activated with high confidence, simply conducting CL or KD without separating co-contexts could bring much noise to the feature representation, consequently limited to suppressing false activation and tackling co-occurrence. Different from previous works, we focus on reducing noise during feature representation from a new perspective of a ’separate and conquer’ paradigm with those techniques, and highlight removing the dependence among co-contexts from two dimensions of image space and feature space, respectively.
@@ 3.1 Problem Definition
The co-occurrence problem stems from the fact that the coupling of co-contexts confuses networks and introduces noise during feature representation, leading to false positive pixels still activated with high confidence. Therefore, the key of SeCo to addressing the problem is to initially separate the co-occurring objects before feature extraction and enhance category-specific representations to promote the discrepancy among co-contexts.
@@ 3.2 Framework Overview
Fig. 2 (a) specifically depicts the pipeline of SeCo. Given a space of input images ${X}$ and a space of classification labels $\mathcal{Y}=\{1,2,\ldots,K\}$, where the number of categories is denoted as $K$, the training dataset is defined as ${D}=$$\left\{\left({I}_{{i}},{Y}_{{i}}\right)\right\}_{{i}=1}^{{V}}$. Each tuple $({I},{Y})$ in ${D}$ is the input, where ${I}\in\mathbb{R}^{3\times H\times W}$ is an image and ${Y}\in\mathcal{Y}$ is the class label. In our framework, we decompose an image $I$ into patches $x$ and remove the spatial dependence among co-categories. Then we build a dual-teacher-single-student architecture to extract multi-granularity knowledge and conduct semantic contrast to facilitate discrepancy of co-contexts.
Fig. 2 (b) illustrates the essence of the proposed losses ${L}_{{LiG}}$, ${L}_{{LiL}}$ and tag rectification strategy. ${L}_{{LiG}}$ means the loss between local patches and global images. The category knowledge (pink circle) from images acts in centroids for the patch representation (colored squares). ${L}_{{LiL}}$ means the loss among local patches. It pushes apart patch semantics in different categories and pulls together those within the same. The rectification strategy excludes noisy patches in abnormal similarity with the help of category tags.
@@ 3.3 Image Decomposition
The essence of contrastive learning is the construction of positive pairs [9]. Simply cropping images into patches cannot help differentiate co-categories [17, 37]. To conduct contrast among co-categories, we assign a category tag $t_{i}$ from the raw CAM seeds to each patch ${x}_{i}$ and guide the subsequent class-specific contrast. Specifically, samples with the same category tags are viewed as positive pairs, while samples with different tags are negative pairs. We firstly generate CAM seeds by introducing an auxiliary classification head in the teacher network. It is found that the auxiliary classification head helps generate diverse CAMs than the final classification head. Although CAM seeds cannot provide precise supervision, this raw signal can effectively guide the initial assignment of category tags. We obtain the auxiliary pseudo mask ${M}_{aux}$ from ${CAM}_{aux}$:
@@
Representation of Local Patches. Following the setups of popular contrastive approaches [15], we generate two augmented views from local patches ${x}$ by implementing weak data augmentation (W.T.) $Aug_{{q}}(\cdot)$ and strong augmentation (S.T.) $Aug_{{k}}(\cdot)$, and send them to the student encoder $g_{q}(\cdot)$ and local teacher encoder $g_{k}(\cdot)$ to extract class embedding ${q_{cls}}$ and ${k_{cls}}$, respectively, as shown in Fig. 2 (a). Class token in ViT is used to represent the embedding as it generalises high-level semantics [4]. We further adopt a MLP operation ${O}_{{q}}(\cdot)$ and ${O}_{{k}}(\cdot)$ on the obtained embeddings to strengthen the feature and obtain the final representation ${q}=\left\{{q}_{{i}}\in\mathbb{R}^{1\times{C}}\right\}_{{i}=1}^{{n}}$ and ${k}=\left\{{k}_{{i}}\in\mathbb{R}^{1\times{C}}\right\}_{{i}=1}^{{n}}$, which denotes the local semantics from patches. The patch representation is formulated as:
@@ 3.4 Representation with Category Knowledge
Class-specific Knowledge Extraction. After spatially separating the co-occurring context, we build a global teacher to dynamically extract category knowledge from integral images. Notably, considering the image decomposition may destruct the semantic context of patches and global CAM lacks local details, we train both teacher and student and share the feature encoders to facilitate the knowledge communication between local and global semantics.
Previous works [8, 28] extract semantics based on CAM. SeCo extracts category knowledge $P$ by utilizing the virtue of class token in ViT [4]. It represents the high-level semantics of each category and avoids the noise from false localization of CAM. In particular, the knowledge set from images consists of $K$ prototypes and each prototype generalises the corresponding category semantics, i.e.,$P=\left\{{P}_{l}\in\mathbb{R}^{1\times C}\right\}_{l=1}^{K}$. Given the input image ${I}$ with categories $l$, we extract the class token $Z_{l}$ from the global teacher encoder $f(\cdot)$ to denote the category representation. The process can be formulated as $Z_{l}=f(I)$.
To reduce the noise from co-occurring objects and comprehensively generalises the knowledge, we propose an adaptive updating strategy to gather all semantics across the dataset. Given the token $Z_{l}$ with multi-class, we calculate the cosine similarity with the corresponding prototypes and leverage the similarity scores after $softmax(\cdot)$ as the weights $W=\left\{{W}_{l}\in\mathbb{R}^{1\times C}\right\}_{l=1}^{K}$ to estimate the relevance to the corresponding category knowledge. The updating process can be formulated as:
@@
Knowledge Guided Contrast for Co-categories. With the global category knowledge and patch semantics $q$, we design ${L}_{LiG}$ loss inspired by InfoNCE [30] to guide the student training. To promote the discrepancy of co-categories, only the co-categories is viewed as negative pairs while the semantics within the same category is positive pairs. Hence, the diversity comparison is held between the filtered local semantics $q_{s}=\{q_{i}\}_{i=1}^{u}$ and the knowledge ${P}_{{s}}$ with the appeared categories, where $u$ is the number of patches. The contrast between patch semantics and category prototypes is achieved by:

### 21 Mar 2024 v3

@@ Abstract
Weakly supervised semantic segmentation (WSSS) with image-level labels aims to achieve segmentation tasks without dense annotations. However, attributed to the frequent coupling of co-occurring objects and the limited supervision from image-level labels, the challenging co-occurrence problem is widely present and leads to false activation of objects in WSSS. In this work, we devise a ’Separate and Conquer’ scheme SeCo to tackle this issue from dimensions of image space and feature space. In the image space, we propose to ’separate’ the co-occurring objects with image decomposition by subdividing images into patches. Importantly, we assign each patch a category tag from Class Activation Maps (CAMs), which spatially helps remove the co-context bias and guide the subsequent representation. In the feature space, we propose to ’conquer’ the false activation by enhancing semantic representation with multi-granularity knowledge contrast. To this end, a dual-teacher-single-student architecture is designed and tag-guided contrast is conducted, which guarantee the correctness of knowledge and further facilitate the discrepancy among co-contexts. We streamline the multi-staged WSSS pipeline end-to-end and tackle this issue without external supervision. Extensive experiments are conducted, validating the efficiency of our method and the superiority over previous single-staged and even multi-staged competitors on PASCAL VOC and MS COCO. Code is available here.
@@ 1 Introduction
Figure 1: (a) Co-occurrence issue. Targets marked by stars (horse and boat) are falsely activated. (b) To solve this issue, we propose a single-staged framework SeCo, which acts in a ’separate and conquer’ manner that efficiently tackles co-occurrence issue without external supervision. It initially separates spatial con-texts in the image space and then conquers false activation in feature space. (c) The proposed SeCo accurately localizes the co-categories.
@@
Formally, the pipeline for WSSS with image-level labels consists of three steps, i.e., initially training a classification model to generate CAM seeds [44], then refining CAMs to generate pseudo labels, and finally training a segmentation network with the pseudo labels and taking it as the final model to inference [2, 47]. Such a WSSS paradigm can be further divided into multi-staged and single-staged groups. For multi-staged WSSS methods [24, 46, 27], the classification and segmentation models need to be trained progressively. It intends to have better segmentation performance while more complicated at training streamline. The single-staged methods [2, 36, 37, 31] share the encoder for classification and segmentation networks, thus can be trained end-to-end. It is more efficient to optimize while holding inferior performance to the multi-staged. In this study, we focus on the most challenging WSSS paradigm with image-level annotations and streamline the paradigm end-to-end.
@@
For both single- and multi-staged WSSS, generating reliable CAMs from image-level labels is the first and fundamental step for the performance [19, 48]. However, since objects intend to co-occur together, such as {train, railroad}, {boat, water}, {horse, person}, etc., it is inevitable to tackle the co-occurrence of objectives in WSSS. The challenging co-occurrence problem is widespread and often leads to false activation [5, 22, 46, 20], as shown in Fig. 1 (a). Although most existing works have succeeded in completing CAMs, they barely pay attention to such issue, consequently limited to tackling the false activation and bottlenecked in WSSS performance. Recently, introducing external supervision or human prior is proposed to tackle this problem.  [46, 27] leverage the vision-language matched CLIP model [33] to help distinguish among coupled contexts. [20] elaborately applies hard out-of-distribution samples to suppress spurious background cues. [39] constructs additional co-categories to address this issue. Although impressive, they heavily rely on external data or elaborate designs to tackle co-occurrence, which impedes real-world applications with complex relations among categories.
@@
Essentially, the co-occurrence problem arises because the co-appeared contexts coupled in images confuse the networks and incur wrong semantic bias during the feature representation, resulting in false positive pixels activated with high probability. Previous methods ignore the importance of separating co-contexts before representation, thus requiring external data or designs to tackle this issue. Based on the analysis above, we argue that initially separating the coupled objects to remove the bias and then enhancing category-specific representation provides a potential insight to address this issue without external supervision.
@@
As illustrated in Fig. 1 (b), we propose a single-staged WSSS framework SeCo that does not require any extra supervision to tackle the problem. Our key insight lies in the ’separate and conquer’ training scheme that decouples co-occurrence in image space and feature space by image decomposition and representation enhancement, respectively. (1) To separate the spatial dependence of co-contexts in image space, we propose to decompose the integral image into multiple patches containing single category information. Previous patch-based method [17] simply gives image-level labels to patches, but the image labels cannot help differentiate the co-contexts at patch level (see Sec. 2.1). Instead, we further explore the strategy to spatially separate coupled objects and focus on tagging each patch. Specifically, category tags from CAMs are designed for each patch, which help identify the co-contexts at patch level. Since CAMs inevitably bring noise in tags, a similarity-based rectification method is designed to revise the noisy tags. In addition, a tag memory pool is constructed to store all history tags, guiding the subsequent patch representation. (2) To conquer the co-context confusion and enhance the semantic representation in feature space, we design a dual-teacher single-student architecture to promote the discrepancy among co-categories. We first build a global teacher to extract category knowledge from integral images. The knowledge provides class centroids for the student in patch representation and helps to push apart co-contexts. Considering the trade-offs between the separation of co-categories and the destruction on global contexts, we share the encoders of both branches to provide complementary information for patch and image semantics. In addition, patch semantic knowledge is further extracted from a patch-level semantic reservoir maintained by a local teacher. Guided by the tags from memory pool, the knowledge helps remove the bias during the representation and pushes apart co-contexts while pulling together those within the same category at a fine-grained level. (3) Along with the category tags and the extracted knowledge, multi-granularity contrast is further proposed across the whole dataset to decouple the co-contexts deeply.
@@
Extensive experiments are conducted on PASCAL VOC and MS COCO, validating its effectiveness in tackling co-occurrence (as shown in Fig. 1 (c)) and the superiority over previous single-staged and multi-staged competitors.
@@ 2 Related Works
Figure 2: (a) Architecture of the proposed SeCo to tackle co-occurrence issue. Specifically, integral images are firstly sent to the global teacher (G-Teacher) to extract the category knowledge and CAMs. Then three types of category tags, i.e., single-category, background and uncertain tags, are generated from CAMs and allocated to patches accordingly. With tags, two views of patches by different augmentations, i.e., weak data augmentation (W.T.) and strong augmentation (S.T.), are sent to the student and local teacher (L-Teacher) branch, respectively. The local teacher stores all the history patches and category tags and generalizes the patch semantic knowledge. Finally, two contrastive losses, ${L}_{{LiG}}$ and ${L}_{{LiL}}$, are conducted to guarantee the decoupling. In addition, CAMs from global teacher are refined as pseudo labels to train the segmentation network. Since the encoders of segmentation model and classification model are shared, our WSSS pipeline can be trained end-to-end. (b) Illustrated essence of the key components in SeCo. More details are introduced in Sec. 3.2
@@ 2.1 Learning from Local Semantics
The image-level labels provide limited supervision to generate high-quality CAMs, which motivates many researches to dig complementary information from local semantics. Decomposing a whole image into local patches offers a practical implementation. L2G [17] crops patches to mine different views of semantics and generates more complete CAMs. PPL [23] utilizes feature patches to explore scattered local details and generates CAMs to cover the whole object. ToCo [37] extracts local semantics from patches and activates more non-discriminative areas. However, simply giving image-level labels to each patch like L2G cannot identify co-categories at patch level, or grouping them into foregrounds and backgrounds as ToCo fails to tell co-categories in the foregrounds. Both of them aggravate the unexpected co-occurring bias during patch representation and suffer from false activation. To this end, we observe that rich semantics from CAMs can be potentially used to tag each patch, which helps differentiate co-contexts and guide category representation at a patch level. To the best of our knowledge, we are the first to report that assigning class-specific tags to patches effectively separates co-categories and reduces the bias during the representation.
@@ 2.2 Contrastive Learning & Knowledge Distillation
Contrastive learning (CL) [9, 15, 43] and knowledge distillation (KD) [16, 4] are two prevalent techniques to promote feature representation. Inspired by it, RCA [52] and ToCo [37] conduct region-level contrast to focus on non-discriminative pixels. PPC [13] leverages prototype-based [51] contrast to expand CAMs. Apart from incorporating CL, SCD [49] and L2G [17] introduce knowledge distillation into WSSS and succeed in completing CAMs. However, since the co-occurring pixels intend to be falsely activated with high confidence, simply conducting CL or KD without separating co-contexts could bring much noise to the feature representation, consequently limited to suppressing false activation and tackling co-occurrence. Different from previous works, we focus on reducing noise during feature representation from a new perspective of a ’separate and conquer’ paradigm with those techniques, and highlight removing the dependence among co-contexts from two dimensions of image space and feature space, respectively.
@@ 3.1 Problem Definition
The co-occurrence problem stems from the fact that the coupling of co-contexts confuses networks and introduces noise during feature representation, leading to false positive pixels still activated with high confidence. Therefore, the key of SeCo to addressing the problem is to initially separate the co-occurring objects before feature extraction and then enhance category-specific representations to promote the discrepancy among co-contexts.
@@ 3.2 Framework Overview
Fig. 2 (a) specifically depicts the pipeline of SeCo. Given a space of input images ${X}$ and a space of classification labels $\mathcal{Y}=\{1,2,\ldots,K\}$, where the number of categories is denoted as $K$, the training dataset is defined as ${D}=$$\left\{\left({I}_{{i}},{Y}_{{i}}\right)\right\}_{{i}=1}^{{V}}$. Each tuple $({I},{Y})$ in ${D}$ is the input, where ${I}\in\mathbb{R}^{3\times H\times W}$ is the image and ${Y}\in\mathcal{Y}$ is the class label. In our framework, we decompose an image $I$ into patches $x$ and remove the spatial dependence among co-categories. Then we build a dual-teacher-single-student architecture to extract multi-granularity knowledge and conduct semantic contrast to facilitate discrepancy among co-contexts.
Fig. 2 (b) illustrates the essence of the proposed losses ${L}_{{LiG}}$, ${L}_{{LiL}}$ and tag rectification strategy. ${L}_{{LiG}}$ means the loss between local patches and global images. The category knowledge (pink circle) from images acts in centroids for the patch representation (colored squares). ${L}_{{LiL}}$ means the loss among local patches. It pushes apart patch semantics in different categories and pulls together those within the same. The rectification strategy excludes noisy patches in abnormal similarity with the help from category tags.
@@ 3.3 Image Decomposition
The essence of contrastive learning is the construction of positive pairs [9]. Simply cropping images into patches cannot help differentiate co-categories [17, 37]. To conduct contrast among co-categories, we assign a category tag $t_{i}$ from the raw CAM seeds to each patch ${x}_{i}$ and leverage tags to guide the subsequent class-specific contrast. Specifically, samples with the same tags are viewed as positive pairs, while samples with the different are negative pairs. We firstly generate CAM seeds by introducing an auxiliary classification head in the teacher network. It is found that the auxiliary head from intermediate features helps generate more diverse CAMs than that from final features [37]. Although CAMs cannot provide precise supervision, this raw signal effectively guides the assignment of category tags. We obtain the auxiliary pseudo mask ${M}_{aux}$ from ${CAM}_{aux}$:
@@
Representation of Local Patches. Following the setups of popular contrastive approaches [15], we generate two augmented views from local patches ${x}$ by implementing weak data augmentation (W.T.) $Aug_{{q}}(\cdot)$ and strong augmentation (S.T.) $Aug_{{k}}(\cdot)$, and send them to the student encoder $g_{q}(\cdot)$ and local teacher encoder $g_{k}(\cdot)$ to extract class embedding ${q_{cls}}$ and ${k_{cls}}$, respectively, as shown in Fig. 2 (a). Class token in ViT is used to represent the embedding as it generalizes high-level semantics [4]. We further adopt a MLP operation ${O}_{{q}}(\cdot)$ and ${O}_{{k}}(\cdot)$ on the obtained embeddings to strengthen the feature and obtain the final representation ${q}=\left\{{q}_{{i}}\in\mathbb{R}^{1\times{C}}\right\}_{{i}=1}^{{n}}$ and ${k}=\left\{{k}_{{i}}\in\mathbb{R}^{1\times{C}}\right\}_{{i}=1}^{{n}}$, which denotes the local semantics from patches. The patch representation is formulated as:
@@ 3.4 Representation with Category Knowledge
Class-specific Knowledge Extraction. After spatially separating the co-occurring context, we build a global teacher to dynamically extract category knowledge from integral images. Notably, considering the image decomposition may destruct the semantic context of patches while global CAMs lack local details, we train both teacher and student and share the encoders to facilitate the knowledge communication between local and global semantics.
Previous works [8, 28] extract semantics based on CAM. SeCo extracts category knowledge $P$ by utilizing the virtue of class token in ViT [4]. It represents the high-level semantics of each category and avoids the noise from false localization of CAMs. In particular, the knowledge set from images consists of $K$ prototypes and each prototype generalizes the corresponding category semantics, i.e.,$P=\left\{{P}_{l}\in\mathbb{R}^{1\times C}\right\}_{l=1}^{K}$. Given the input image ${I}$ with categories $l$, we extract the class token $Z_{l}$ from the global teacher encoder $f(\cdot)$ to denote the category representation. The process can be formulated as $Z_{l}=f(I)$.
To reduce the noise from co-occurring objects and comprehensively generalize the knowledge, we propose an adaptive updating strategy to gather all semantics across the dataset. Given the token $Z_{l}$ with multi-class, we calculate the cosine similarity with the corresponding prototypes and leverage the similarity scores after $softmax(\cdot)$ as the weights $W=\left\{{W}_{l}\in\mathbb{R}^{1\times C}\right\}_{l=1}^{K}$ to estimate the relevance to the corresponding category knowledge. The updating process can be formulated as:
@@
Knowledge Guided Contrast for Co-categories. With the global category knowledge and patch semantics $q$, we design ${L}_{LiG}$ loss inspired by InfoNCE [30] to guide the student training. To promote the discrepancy of co-categories, only the co-categories are viewed as negative pairs while the semantics within the same category is positive pairs. Hence, the diversity comparison is held between the filtered local semantics $q_{s}=\{q_{i}\}_{i=1}^{u}$ and the knowledge ${P}_{{s}}$ with the appeared categories, where $u$ is the number of patches. The contrast between patch semantics and category prototypes is achieved by: